OpenStudy (anonymous):

A cannonball when configured to be fired at a certain angle would have a parabolic path with a maximum height of 100m and a horizontal range of 40m. if the cannonball is placed at the edge of the cliff 125 high, how far from the base would the cannon ball land?

OpenStudy (jack1):

\(\large s = u \times t + \frac 12 \times a \times t^2\) u= initial velocity t = time a = acceleration this is one of the kinematic equations... there are a few, does it look like something you've seen before? you can use this equation and the others, in both the horizontal and vertical planes, to work out the final range of the cannonball

OpenStudy (jack1):